### v1

# Runyang You1    Yongqi Li1   Xinyu Lin2   Xin Zhang1 Wenjie Wang3   Wenjie Li1   Liqiang Nie4 1

Abstract
Large recommender models have extended LLMs as powerful recommenders via encoding or item generation, and recent breakthroughs in LLM reasoning synchronously motivate the exploration of reasoning in recommendation. Current studies usually position LLMs as external reasoning modules to yield auxiliary thought for augmenting conventional recommendation pipelines. However, such decoupled designs are limited in significant resource cost and suboptimal joint optimization. To address these issues, we propose $\text{R}^{2}\text{ec}$, a unified large recommender model with intrinsic reasoning capabilities. Initially, we reconceptualize the model architecture to facilitate interleaved reasoning and recommendation in the autoregressive process. Subsequently, we propose RecPO, a corresponding reinforcement learning framework that optimizes $\text{R}^{2}\text{ec}$ both the reasoning and recommendation capabilities simultaneously in a single policy update; RecPO introduces a fused reward scheme that solely leverages recommendation labels to simulate the reasoning capability, eliminating dependency on specialized reasoning annotations. Experiments on three datasets with various baselines verify the effectiveness of $\text{R}^{2}\text{ec}$, showing relative improvements of 68.67% in Hit@5 and 45.21% in NDCG@20. Code available at https://github.com/YRYangang/RRec.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in contextual understanding and open-ended generation [1, 2, 3, 4]. This success has catalyzed the development of large recommender models that inherit and specialize in the advantages of LLMs for recommendation. Current approaches can be divided into two main streams: one employs LLMs as powerful encoders to embed users (their historical interactions) [5, 6, 7], while the other reformulates item prediction into the autoregressive generation of item identifiers [4, 8]. These large recommender models exhibit remarkable generalization capabilities, achieving advanced performance across diverse application scenarios, including cold-start recommendations [9], cross-domain personalization [10], and long-tail item prediction [11, 12].
The frontier of LLM capabilities now extends beyond model size scaling to test-time scaling, i.e., reasoning. Emerging advances like DeepSeek-R1 [13] demonstrates that such extra computation during inference can further improve LLM capabilities in areas such as mathematics, coding, and scientific problem-solving [14]. Given that large recommender models are instantiated from pretrained LLMs, a natural question is: how can large recommender model benefit from reasoning to further enhance recommendation performance?
Existing studies have preliminarily explored LLM reasoning for recommendation, including user preference analysis [15, 16], synthetic item profile enrichment [17, 18], user search query rewriting [19, 20], and rating prediction [21]. These approaches typically take LLMs as additional reasoning modules to augment the original recommendation pipeline, decoupling reasoning from item prediction. However, such designs introduce crucial limitations: 1) Significant resource cost. Two distinct modules, namely the large reasoning model and the recommendation model, must be trained, checkpointed, and served in parallel, inflating both memory footprint and inference latency. 2) Suboptimal joint optimization. Reasoning and recommendation modules can only be trained by freezing one while updating the other—first update the reasoning module, then switch to the recommendation module, and continue alternating in this manner [15]. This alternate optimization scheme prevents gradient flow across the pipeline, precluding true end-to-end learning, hindering fine-grained alignment between reasoning and ranking objectives, and ultimately leading to suboptimal convergence. In this work, we aim to develop a unified large recommender model with intrinsic reasoning capabilities, exploring new technical solutions for reasoning-enhanced recommendation systems.
Figure 1: The architecture overview of $\text{R}^{2}\text{ec}$, which facilitates interleaved reasoning and recommendation in an autoregressive process with two task-specific heads: 1) language-modeling head (lm_head) for reasoning generation; and 2) recommendation head (rec_head) for item prediction.
It is non-trivial to develop a unified large recommender model with reasoning due to the following aspects: 1) Model design. Recommendation is traditionally framed as a discriminative ranking task, whereas reasoning is a generative process. How to incorporate both generative reasoning and discriminative recommendation into one single model architecture is worth careful consideration. 2) Training optimization. It is also challenging to stimulate and optimize the reasoning capabilities of large recommender models. On the one hand, recommendation domains lack annotated reasoning traces, as personalized recommendation rationale is hardly obtainable, making supervised learning impractical. On the other hand, Reinforcement Learning (RL) [20, 22, 23] is a natural alternative but encounters unique problems in reward and training objective design within the recommendation task. Since reasoning and recommendation capabilities interact within a single model, this necessitates a training framework that effectively bridges these two aspects.
To address the above issues, we propose a unified large Recommender model with intrinsic Reasoning, dubbed as $\text{R}^{2}\text{ec}$. 1) In the model design, as illustrated in Figure 1, we introduce the recommendation head to the LLM backbone, alongside the original language-modeling head. This setup allows the model to first generate reasoning tokens autoregressively and then predict the item in the last step. This model architecture integrates both generation and recommendation tasks, while the next crucial step is to simulate the reasoning capabilities tailored for accurate item predictions. 2) In training optimization, we propose an RL-based training framework without human reasoning annotations. First, we sample trajectories from the model, which essentially consist of reasoning-then-item sequences, similar to the reasoning-then-answer sequences in LLMs. Second, unlike Deepseek-R1, which defines rewards based on answer accuracy, we found it insufficient to set the reward solely using item recommendation metrics. Therefore, we propose a fused reward scheme that combines discrete ranking rewards with continuous similarity rewards. With the sampled trajectories and assigned rewards, a joint reasoning-and-recommendation training objective is introduced to optimize the model for improving the recommendation through reasoning.
We conducted extensive experiments to compare $\text{R}^{2}\text{ec}$ with various traditional, LLM-based, and reasoning-augmented recommendation baselines. The experimental results on three datasets illustrate $\text{R}^{2}\text{ec}$ significantly outperforms all baselines with relative improvements of 68.67% in Hit@5 and 45.21% in NDCG@20, verifying the effectiveness of the unified reasoning-recommendation model design. We also performed the ablation study and 5 in-depth analysis experiments, demonstrating that $\text{R}^{2}\text{ec}$ possesses strong generalization capabilities with different optimization algorithms. More surprisingly, we observed a similar phenomenon in LLM reasoning that the reasoning length increases as training progresses.
The key contributions in this work are summarized as follows:
{adjustwidth}
[…]
2 Related Work
2.1 Reinforcement Learning for LLM Reasoning
OpenAI’s GPT-o1 demonstrated that scaling reinforcement learning (RL) training with large compute budgets enables models to internalize reasoning as a differentiable policy, achieving state-of-the-art performance in emergent meta-reasoning tasks [24]. Early RLHF methods trained a reward model on human preference data and fine-tuned the policy using Proximal Policy Optimization (PPO) [23]; however, PPO’s multiple optimization passes introduce implementation complexity. To streamline RL tuning, Direct Preference Optimization (DPO) [25] performs single-step policy updates, trading simplicity for potential off-policy bias. Alternative estimators, such as Group Relative Policy Optimization (GRPO) [22], derive baselines from group-level rewards, eliminating the need for a separate critic, while REINFORCE’s leave-one-out (RLOO) [26] computes advantages by excluding each trajectory sequentially. Building on these advancements, DeepSeek-Zero [27] removes reliance on supervised fine-tuning by driving reasoning emergence purely through intrinsic RL rewards, and DeepSeek-R1 further integrates rule-driven reward engineering with self-play verification to enhance reasoning robustness.
2.2 Large Language Model-based Recommendation
Recent advances in LLM-based recommendation typically follow two paradigms: 1) LLM-enhanced, where LLMs enrich recommendation pipelines with additional features, and 2) LLM-centric, where recommendation is framed as a generative task via item identifiers [1, 2, 3, 4, 28, 29], or LLMs serve as encoders to embed users and items [6, 30]. We term models under the second paradigm as large recommender models. In the first paradigm, LLM-generated embeddings are utilized to enhance conventional recommenders. Adoption of LLM-generated text are not new, typical examples includes user intent [15, 31], search queries [20], item summaries [18], or rationales [32]. Among them, SLIM [32], LangPTune [15], and Rec-R1 [20] further optimize LLM for finer rationale generation with RL. However, such designs introduce high resource demands for training and serving both models, and the connection between two modules remains gradient-less, blocking true end-to-end learning, resulting in suboptimal convergence and performance. Large recommender models typically recommend through autoregressive generation of item identifiers [4, 28] or titles [1, 2, 3, 29], or by embedding user interaction sequences [6, 30]. Large recommender model has demonstrated remarkable potential, but current methods fail to utilize reasoning ability of LLMs. To the best of our knowledge, no existing large recommender model improves recommendation over explicit reasoning generation, which is the gap that motivated our work.
3 Preliminaries
3.1 LLM-based Recommendation
LLM-based recommendation is an emerging paradigm, and the most applicable approach is to leverage LLMs as encoders to embed users and items, which are widely adopted in industry [30, 28, 33]. Typically, the recommendation process within this paradigm involves the following structured steps:
User and Item Formulation.
Structured interaction histories and item metadata are first formulated into natural language prompts suitable for LLMs. We first collect a user’s historically interacted items with the corresponding ratings. We then construct a textual prompt includes both the instruction and a natural-language description of these past behaviors. E.g., User has purchased and rated the following products sequentially: 1. ‘Avatar Blu-ray’ (4/5); 2. ‘Wireless Headphones’ (5/5); … Recommend the next item.” Finally, we tokenize this prompt into a sequence of tokens, denoted as ${x}_{u}$, which serves as the LLM input. Likewise, each candidate item $v$ can be described using its metadata, then tokenized into a sequence of tokens ${x}_{v}$.
The tokenized textual prompts ${x}_{u}$ and ${x}_{v}$ are then input into the LLM transformer backbone, by extracting the final hidden state of the input sequence, we can obtain corresponding user embedding $\bm{h}_{u}$ and item embedding $\bm{h}_{v}$.
Training via Contrastive Learning.
To optimize LLM for recommendation tasks, in-batch contrastive loss is typically adopted:
$\mathcal{L}_{\mathrm{CL}}=-\log\frac{\exp\bigl{(}\bm{h}_{u}^{\top}\bm{h}_{v^{+}}/\tau\bigr{)}}{\sum_{v^{\prime}\in B}\exp\bigl{(}\bm{h}_{u}^{\top}\bm{h}_{v^{\prime}}/\tau\bigr{)}},$ (1)
where $\tau$ is the temperature hyperparameter controlling similarity dispersion, $B$ is the ground truth item set in one batch, and $v^{+}$ is the groud-truth target item.
Inference.
During inference, recommendation score $s(u,v)$ for each user-item pair is computed via the inner product between their embeddings: $s(u,v)=\bm{h}_{u}^{\top}\bm{h}_{v}.$ Finally, we rank all items according to the scores and obtain top-$K$ item recommendation list.
3.2 Reinforcement Learning for Large Reasoning Models
Terminologies and Notations.
RL for LLMs treats text generation as a sequential action process, where each action refers to outputting the next token. Formally, at each time step $t$, we denote the current state as $(x,o_{<t})$, where $x$ is the initial prompt and $o_{<t}$ is all previously generated tokens. Then, the conditioned probability distribution of an action is denoted as $\pi_{\theta}(o_{t}|x,o_{<t})$, where $\pi_{\theta}$ is the policy parameterized by $\theta$. Based on the distribution, we can obtain the next token $o_{t}$ through manual selection, greedy or temperature sampling. Upon obtaining a complete sequence $o$, termed as a trajectory, a reward function then evaluates and assigns it with a scalar score $R$ that can reflect user satisfaction or answer correctness. However, using $R$ directly in gradient estimates leads to high variance and unstable updates; instead, the advantage $A$, which measures how much better an action is than the expected return from that state, is adopted to reduce variance and improve sample efficiency [22, 23, 26, 34].
Sampling and Advantage Computation.
Sampling-based advantage estimation is widely adopted in recent advances [22, 26, 34, 35]; below we describe its basic pipeline and two typical computation methods. Given an input ${x}$, a group of $G$ different trajectories $\{{o}_{i}\}_{i=1}^{G}$, are sampled from $\pi_{\theta_{\text{old}}}$. Existing studies widely obtain these trajectories via top-$K$ sampling with temperature, where $\theta_{\text{old}}$ refers to the frozen policy parameters before current update. Each trajectory $o_{i}$ receives a scalar reward $R_{i}$, which will then be used to compute the trajectory-level advantages via two widely adopted approaches, namely GRPO [22] and RLOO [26]:
$A^{\mathrm{RLOO}}_{i}=R_{i}-\frac{1}{G-1}\sum_{j\neq i}^{G}R_{j},\quad A^{\mathrm{GRPO}}_{i}=\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{G})}{\mathrm{std}(\{R_{j}\}_{j=1}^{G})}.$ (2)
Training Objective.
Training then proceeds with policy-gradient algorithms. Specifically, let
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|x,{o}_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|x,{o}_{i,<t})}$ (3)
be the importance ratio between the updated and old policies of trajectory $i$ at token position $t$. The training objective is given by:
$\mathcal{J}(\theta)=\mathbb{E}_{{x}\sim\mathcal{D},\{{o}_{i}\}^{G}_{i=1}\sim\pi_{\theta_{\text{old}}}(\cdot|{x})}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left[\min\left(r_{i,t}(\theta)A_{i},\ \text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)\right],$ (4)
where $\epsilon$ defines the clipping trust-region radius, which prevents excessively large updates, thereby reducing variance and improving optimization stability.
4 $\text{R}^{2}\text{ec}$: Large Recommender Model with Reasoning
$\text{R}^{2}\text{ec}$ is a large recommender model that “thinks” to recommend. We first outline the model design that enables the incorporation of generative reasoning and discriminative recommendation into a single model in Section 4.1. And then our proposed RL optimization algorithm tailored for achieving unified reasoning and recommendation capabilities is introduced in Section 4.2.
4.1 Model Design
Architecture.
As illustrated in Figure 1, our proposed $\text{R}^{2}\text{ec}$ is built upon a decoder-only backbone with two task-specific heads: 1) Language-modeling head (lm_head). The token embedding table $\mathbf{H}_{\mathcal{T}}\in\mathbb{R}^{|\mathcal{T}|\times d}$. Where $\mathcal{T}$ is the token set, and each row is a $d$-dimensional embedding for one token. This head is responsible for the generation of reasoning tokens. 2) Recommendation head (rec_head). The item embedding table $\mathbf{H}_{\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times d}$. Where $\mathcal{V}$ is the item set, and each row $\bm{h}_{v}$ in $\mathbf{H}_{\mathcal{V}}$ is obtained by encoding item description prompt into the model itself and extracting the final hidden state. This head is used to score items for recommendation. Such design integrates generative reasoning and discriminative recommendation within a single unified model. It supports flexible and scalable item update by simply adding, deleting, or replacing vectors in the item embedding table, contrasting with generative recommendation systems that require hard-coded tokenization  [36, 8, 4, 37], enabling effective zero-shot generalization and accommodating large-scale item catalogs without severe degradation on recommendation quality or efficiency.
Item prediction.
Inference begins with feeding the tokenized user prompt ${x}_{u}$ (template of prompts can be found in Appendix E) into the transformer-based backbone, producing an initial hidden state $\bm{h}_{0}$. The language-modeling head then maps $\bm{h}_{0}$ to the first reasoning token $o_{1}$. This process continues autoregressively, yielding a sequence of $T$ reasoning tokens ${o}_{1:T}$. The final hidden state of the generated sequence $\bm{h}_{T}$ is then fed into the recommendation head, where each candidate item $v\in\mathcal{V}$ is scored by an inner product $s(v)={\bm{h}}_{T}^{\top}\mathbf{H}_{\mathcal{V}}[v],$ $v\in\mathcal{V},$ which determines the final ranking. This mechanism yields a tight reasoning–recommendation coupling, since both the language-modeling head and recommendation head share the same hidden-state space as input, reasoning directly reshapes $\bm{h}_{T}$ and thus yielding more accurate recommendation scores $s(v)$. Such alignment ensures that reasoning optimization (Section 4.2) contributes directly to finer recommendation.
4.2 Optimization: RecPO
Our goal is to train the policy $\pi_{\theta}$ to jointly performs reasoning and recommendation, i.e., it must generate coherent reasoning sequences to rank the target item accurately. Accordingly, we structure our optimization workflow in three parts. First, we introduce the trajectory sampling strategy that draws multiple reasoning trajectories for each user (Section 4.2.1). Second, we describe reward and advantage estimation, where discrete ranking signals and softmax similarities are fused into a single scalar award for the above sampled trajectories (Section 4.2.2). Finally, we formulate the training objective, which blends reasoning-level and recommendation-level updates through a clipped-ratio loss (Section 4.2.3). For a complete description of our end-to-end training and inference pipeline, please refer to Appendix D.
4.2.1 Trajectory Sampling
Due to the optimization objective of joint learning, we define one trajectory in our settings spanning the entire reasoning-then-recommend process:
$x_{u}\;\xrightarrow{\;\pi_{\theta}\;}\;o_{1}\;\xrightarrow{\;\pi_{\theta}\;}\;...\;\xrightarrow{\;\pi_{\theta}\;}\;o_{T}\;\xrightarrow{\;\pi_{\theta}\;}\;v^{+},$
where the initial state $x_{u}$ encodes the user history and instruction, $o_{1},...,o_{T}$ represents the intermediate actions of outputing $T$ reasoning tokens, and $v^{+}$ as the action of recommending the ground-truth target item.
For each user $u$, we first sample $G$ reasonings $\{{o}_{i}\}_{i=1}^{G}$ with the old policy $\pi_{\theta_{\text{old}}}$ using the tokenized input $x_{u}$: $\{{o}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x_{u})$ by top-$K$ sampling with temperature to control stochasticity. Each sampled reasoning is then fed through the policy $\pi_{\theta}$ to produce a complete reasoning-then-recommend trace, which is subsequently used for reward calculation and advantage estimation.
4.2.2 Reward and Advantage Estimation
Given the above sampled trajectories, we now aim to assign rewards to them. Basically, the reward should align with the evaluation criteria, i.e., the recommendation metrics in our work, encouraging the model to achieve better performance. However, in practice, we find that directly using the recommendation metrics as rewards is insufficient, as many trajectories of varying quality can result in the same top-$K$ ranking. We therefore introduce a fused reward scheme that combines discrete ranking reward $R_{d}$ and continuous similarity reward $R_{c}$, which are formulated as follows:
${R_{d}}=\operatorname{NDCG}@k\left(\text{rank}(v^{+})\right),\qquad{R_{c}}=\frac{\exp\bigl{(}\bm{h}_{T}^{\top}\bm{h}_{v^{+}}/\tau\bigr{)}}{\sum_{v\in\mathcal{V}}\exp\bigl{(}\bm{h}_{T}^{\top}\bm{h}_{v}/\tau\bigr{)}},$
where $R_{d}$ is the NDGC, $R_{c}$ is the softmax similarity of recommending target item against all items in $\mathcal{V}$. The final reward is then obtained through a linear combination:
$R=\beta\,R_{c}+(1-\beta)\,R_{d},\qquad\beta\in[0,1],$ (5)
where the weighting coefficient $\beta$ is empirically set to $\beta\approx 0.05$ to keep the ranking term dominant while providing sufficient resolution among trajectories that attain identical ranks. With rewards $\{R_{i}\}_{i=1}^{G}$ we can obtain trajectory-level advantages $\{A_{i}\}_{i=1}^{G}$ via Eqn.(2).
4.2.3 Training Objective
Given the goal of joint optimization of reasoning and recommendation, we treat the entire reasoning-then-recommend sequence $(x_{u},o_{1},...,o_{T},v^{+})$ as a single RL trajectory. Policy optimization therefore operates over a composite action space, where the policy first makes token-level decisions to generate reasoning, then selects an item at the recommendation stage. Under this formulation, the importance ratio from Eqn. (3) extends to:
$\quad r_{i,t}(\theta)$ =$=\begin{cases}\frac{\pi_{\theta}(o_{i,t}|x_{u},{o}_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|x_{u},{o}_{i,<t})},&\text{if }t\leqslant T\text{ (reasoning)}\\ \frac{\pi_{\theta}(v^{+}|x_{u},{o}_{i,\leqslant T})}{\pi_{\theta_{\text{old}}}(v^{+}|x_{u},{o}_{i,\leqslant T})},&\text{if }t=T+1\text{ (recommendation)}.\end{cases}$ (6)
Specifically, we model the recommending action, i.e., recommending the target item $v^{+}$ via the in-batch softmax:
$\pi_{\theta}(v^{+}\,|\,{x}_{u},{o}_{i})=\frac{\exp\bigl{(}\bm{h}_{T}^{\top}\bm{h}_{v^{+}}/\tau\bigr{)}}{\sum_{v^{\prime}\in{B}}\exp\bigl{(}\bm{h}_{T}^{\top}\bm{h}_{v^{\prime}}/\tau\bigr{)}}.$ (7)
Let $\ell_{\epsilon}(r,A)=\min(rA,\;\operatorname{clip}(r,1-\epsilon,1+\epsilon)A)$ be the standard clipping operator with threshold $\epsilon$, we define the joint reasoning-and-recommendation training objective as:
$\mathcal{J}(\theta)=\mathbb{E}_{\{u,v^{+}\}\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x_{u})}\frac{1}{G}\sum_{i=1}^{G}\Bigl{[}\sum_{t=1}^{T_{i}}\ell_{\epsilon}(r_{i,t}(\theta),A_{i})+\delta_{i,i^{\star}}\ell_{\epsilon}\!\bigl{(}r_{i,T+1}(\theta),A_{i}\bigr{)}\Bigr{]}.$ (8)
In Eqn.(8), all trajectories contribute to token-level policy updates via the standard clipped objective $\ell_{\epsilon}\!\bigl{(}r_{i,t}(\theta),A_{i}\bigr{)}$. This design ensures the policy continues to learn from diverse reasoning behaviors. For the last recommendation action, only the trajectory with the highest advantage, identified by $i^{\star}=\arg\max_{j}A_{j}$, contributes gradients to recommendation optimization. This design concentrates recommendation learning with the most promising reasoning path, while the presence of other $G-1$ reasoning paths contributes to ample exploration in reasoning actions, thus preserving exploration and ensuring effective recommendation learning.
5 Experiments
5.1 Setups
Dataset and Metrics.
Following previous work on large recommender models [3, 2, 1, 29], we conducted experiments using real-world datasets sourced from Amazon, including CDs and Vinyl (CDs), Video Games (Games), and Musical Instruments (Instruments). Dataset statistics and preprocessing steps can be referred to Appendix A. We utilized two commonly used metrics: Hit Rate (H@K) and Normalized Discounted Cumulative Gain (N@K), with cutoff $K$ set to $5$, $10$ and $20$. We adopted the full set recommendation setting, where metrics are computed over the entire item set, providing a better reflection of practical scenarios.
Baselines and Implementation.
To comprehensively evaluate $\text{R}^{2}\text{ec}$, we selected competitive baselines from various categories, including traditional sequential recommenders (GRU4Rec [38], Caser [39], SASRec [40]), LLM-based recommender (TIGER [8], BigRec [2], $D^{3}$ [3]), and reasoning augmented recommendation systems (LangPTune [15]). More baseline details can be found in Appendix B. We used Qwen and Gemma as the LLM backbones. More implementation details are described in Appendix C.
5.2 Overall Performance
Table 1: The overall performance of baselines and $\text{R}^{2}\text{ec}$ on three datasets. The best results in each group are marked in Bold, while the second-best results are underlined. * implies the improvements over the second-best results are statistically significant (p-value < 0.05). % improve represents the relative improvement achieved by $\text{R}^{2}\text{ec}$ over the best baseline.
Instruments CDs and Vinyl Video Games
Method H@5 N@5 H@10 N@10 H@20 N@20 H@5 N@5 H@10 N@10 H@20 N@20 H@5 N@5 H@10 N@10 H@20 N@20
GRU4Rec 0.017 067 8340.0170678340.017\,067\,8340.017 067 834 0.013 454 7490.0134547490.013\,454\,7490.013 454 749 0.019 256 0180.0192560180.019\,256\,0180.019 256 018 0.014 156 5580.0141565580.014\,156\,5580.014 156 558 0.020 131 2910.0201312910.020\,131\,2910.020 131 291 0.014 363 2630.0143632630.014\,363\,2630.014 363 263 0.006 6960.0066960.006\,6960.006 696 0.003 6820.0036820.003\,6820.003 682 0.010 4170.0104170.010\,4170.010 417 0.004 1410.0041410.004\,1410.004 141 0.015 6250.0156250.015\,6250.015 625 0.005 0690.0050690.005\,0690.005 069 0.010 8640.0108640.010\,8640.010 864 0.006 9500.0069500.006\,9500.006 950 0.018 1070.0181070.018\,1070.018 107 0.009 2950.0092950.009\,2950.009 295 0.030 0740.0300740.030\,0740.030 074 0.012 3130.0123130.012\,3130.012 313
Caser 0.010 905 4080.0109054080.010\,905\,4080.010 905 408 0.014 053 9050.0140539050.014\,053\,9050.014 053 905 0.011 452 3670.0114523670.011\,452\,3670.011 452 367 0.014 912 6430.0149126430.014\,912\,6430.014 912 643 0.012 660 7920.0126607920.012\,660\,7920.012 660 792 0.015 461 320.015461320.015\,461\,320.015 461 32 0.004 464 2860.0044642860.004\,464\,2860.004 464 286 0.002 918 8690.0029188690.002\,918\,8690.002 918 869 0.006 696 4290.0066964290.006\,696\,4290.006 696 429 0.003 679 9360.0036799360.003\,679\,9360.003 679 936 0.008 928 5710.0089285710.008\,928\,5710.008 928 571 0.004 241 4690.0042414690.004\,241\,4690.004 241 469 0.012 4390.0124390.012\,4390.012 439 0.008 2630.0082630.008\,2630.008 263 0.019 0520.0190520.019\,0520.019 052 0.010 3470.0103470.010\,3470.010 347 0.027 8700.0278700.027\,8700.027 870 0.012 5740.0125740.012\,5740.012 574
SASRec 0.017 505 470.017505470.017\,505\,470.017 505 47 0.014 442 0130.0144420130.014\,442\,0130.014 442 013 0.020 131 2910.0201312910.020\,131\,2910.020 131 291 0.016 192 560.016192560.016\,192\,560.016 192 56 0.022 319 4750.0223194750.022\,319\,4750.022 319 475 0.021 006 5650.0210065650.021\,006\,5650.021 006 565 0.007 583 7320.0075837320.007\,583\,7320.007 583 732 0.010 416 6670.0104166670.010\,416\,6670.010 416 667 0.008 079 7640.0080797640.008\,079\,7640.008 079 764 0.011 904 7620.0119047620.011\,904\,7620.011 904 762 0.008 635 140.008635140.008\,635\,140.008 635 14 0.014 136 9050.0141369050.014\,136\,9050.014 136 905 0.012 9110.0129110.012\,9110.012 911 0.007 9860.0079860.007\,9860.007 986 0.020 6270.0206270.020\,6270.020 627 0.010 4510.0104510.010\,4510.010 451 0.032 5930.0325930.032\,5930.032 593 0.013 4650.0134650.013\,4650.013 465
TIGER 0.017 067 8340.0170678340.017\,067\,8340.017 067 834 0.012 779 1160.0127791160.012\,779\,1160.012 779 116 0.018 380 7440.0183807440.018\,380\,7440.018 380 744 0.013 207 3890.0132073890.013\,207\,3890.013 207 389 0.019 256 0180.0192560180.019\,256\,0180.019 256 018 0.013 436 5330.0134365330.013\,436\,5330.013 436 533 0.006 696 4290.0066964290.006\,696\,4290.006 696 429 0.004 490 7480.0044907480.004\,490\,7480.004 490 748 0.009 672 6190.0096726190.009\,672\,6190.009 672 619 0.005 468 7790.0054687790.005\,468\,7790.005 468 779 0.015 6250.0156250.015\,6250.015 625 0.006 920 2180.0069202180.006\,920\,2180.006 920 218 0.012 2820.0122820.012\,2820.012 282 0.008 4540.0084540.008\,4540.008 454 0.022 2010.0222010.022\,2010.022 201 0.011 6290.0116290.011\,6290.011 629 0.032 2780.0322780.032\,2780.032 278 0.014 1890.0141890.014\,1890.014 189
Qwen BigRec 0.005 204 944 697 462 590.005204944697462590.005\,204\,944\,697\,462\,590.005 204 944 697 462 59 0.003 322 852 180 568 223 40.00332285218056822340.003\,322\,852\,180\,568\,223\,40.003 322 852 180 568 223 4 0.011 060 507 482 108 0020.0110605074821080020.011\,060\,507\,482\,108\,0020.011 060 507 482 108 002 0.005 186 569 480 914 7830.0051865694809147830.005\,186\,569\,480\,914\,7830.005 186 569 480 914 783 0.018 867 924 528 301 8860.0188679245283018860.018\,867\,924\,528\,301\,8860.018 867 924 528 301 886 0.007 178 356 002 302 60.00717835600230260.007\,178\,356\,002\,302\,60.007 178 356 002 302 6 0.004 464 285 714 285 7140.0044642857142857140.004\,464\,285\,714\,285\,7140.004 464 285 714 285 714 0.002 468 399 825 377 9280.0024683998253779280.002\,468\,399\,825\,377\,9280.002 468 399 825 377 928 0.008 928 571 428 571 4280.0089285714285714280.008\,928\,571\,428\,571\,4280.008 928 571 428 571 428 0.003 868 539 672 861 917 50.00386853967286191750.003\,868\,539\,672\,861\,917\,50.003 868 539 672 861 917 5 0.014 136 904 761 904 7620.0141369047619047620.014\,136\,904\,761\,904\,7620.014 136 904 761 904 762 0.005 224 595 192 621 3440.0052245951926213440.005\,224\,595\,192\,621\,3440.005 224 595 192 621 344 0.000 787 640.000787640.000\,787\,640.000 787 64 0.000 388 510.000388510.000\,388\,510.000 388 51 0.001 575 20.00157520.001\,575\,20.001 575 2 0.000 641 110.000641110.000\,641\,110.000 641 11 0.012 759 9240.0127599240.012\,759\,9240.012 759 924 0.003 378 40.00337840.003\,378\,40.003 378 4
$D^{3}$ 0.004 231 770 833 333 3330.0042317708333333330.004\,231\,770\,833\,333\,3330.004 231 770 833 333 333 0.001 951 780 712 180 1950.0019517807121801950.001\,951\,780\,712\,180\,1950.001 951 780 712 180 195 0.009 440 104 166 666 6660.0094401041666666660.009\,440\,104\,166\,666\,6660.009 440 104 166 666 666 0.003 726 565 934 774 1340.0037265659347741340.003\,726\,565\,934\,774\,1340.003 726 565 934 774 134 0.019 205 729 166 666 6680.0192057291666666680.019\,205\,729\,166\,666\,6680.019 205 729 166 666 668 0.006 168 169 982 327 4740.0061681699823274740.006\,168\,169\,982\,327\,4740.006 168 169 982 327 474 0.008 184 523 809 523 810.008184523809523810.008\,184\,523\,809\,523\,810.008 184 523 809 523 81 0.005 685 264 273 880 729 50.00568526427388072950.005\,685\,264\,273\,880\,729\,50.005 685 264 273 880 729 5 0.014 136 904 761 904 7620.0141369047619047620.014\,136\,904\,761\,904\,7620.014 136 904 761 904 762 0.007 614 619 350 554 6390.0076146193505546390.007\,614\,619\,350\,554\,6390.007 614 619 350 554 639 0.025 297 619 047 619 0480.0252976190476190480.025\,297\,619\,047\,619\,0480.025 297 619 047 619 048 0.010 435 130 188 756 8430.0104351301887568430.010\,435\,130\,188\,756\,8430.010 435 130 188 756 843 0.005 356 020.005356020.005\,356\,020.005 356 02 0.002 8010.0028010.002\,8010.002 801 0.010 396 9750.0103969750.010\,396\,9750.010 396 975 0.004 394 20.00439420.004\,394\,20.004 394 2 0.019 691 240.019691240.019\,691\,240.019 691 24 0.006 707 90.00670790.006\,707\,90.006 707 9
LangPTune 0.012 6870.0126870.012\,6870.012 687 0.008 342 80.00834280.008\,342\,80.008 342 8 0.022 4460.0224460.022\,4460.022 446 0.011 4640.0114640.011\,4640.011 464 0.034 8080.0348080.034\,8080.034 808 0.014 5080.0145080.014\,5080.014 508 0.007 440 50.00744050.007\,440\,50.007 440 5 0.005 332 20.00533220.005\,332\,20.005 332 2 0.015 6250.0156250.015\,6250.015 625 0.007 990 60.00799060.007\,990\,60.007 990 6 0.020 8330.0208330.020\,8330.020 833 0.009 404 90.00940490.009\,404\,90.009 404 9 0.004 881 121 083 293 9690.0048811210832939690.004\,881\,121\,083\,293\,9690.004 881 121 083 293 969 0.002 728 30.00272830.002\,728\,30.002 728 3 0.008 815 70.00881570.008\,815\,70.008 815 7 0.003 979 20.00397920.003\,979\,20.003 979 2 0.014 0140.0140140.014\,0140.014 014 0.014 0140.0140140.014\,0140.014 014
$\text{R}^{2}\text{ec}$ 0.023 7480.0237480.023\,7480.023 748* 0.015 3850.0153850.015\,3850.015 385* 0.037 410 540.037410540.037\,410\,540.037 410 54* 0.019 7870.0197870.019\,7870.019 787* 0.061 4830.0614830.061\,4830.061 483* 0.025 8510.0258510.025\,8510.025 851* 0.051 3390.0513390.051\,3390.051 339* 0.037 1660.0371660.037\,1660.037 166* 0.064 7320.0647320.064\,7320.064 732* 0.041 3890.0413890.041\,3890.041 389* $0.0818$* 0.045 7480.0457480.045\,7480.045 748* 0.028 814 359 943 316 0130.0288143599433160130.028\,814\,359\,943\,316\,0130.028 814 359 943 316 013* 0.018 453 733 238 004 8830.0184537332380048830.018\,453\,733\,238\,004\,8830.018 453 733 238 004 883* 0.053 219 965 359 785 860.053219965359785860.053\,219\,965\,359\,785\,860.053 219 965 359 785 86* 0.026 357 850 255 487 170.026357850255487170.026\,357\,850\,255\,487\,170.026 357 850 255 487 17* 0.082 664 147 378 365 610.082664147378365610.082\,664\,147\,378\,365\,610.082 664 147 378 365 61* 0.033 710 982 335 531 660.033710982335531660.033\,710\,982\,335\,531\,660.033 710 982 335 531 66*
% Improve. 35.428 571 428 571 4135.4285714285714135.428\,571\,428\,571\,4135.428 571 428 571 41% 6.944 444 444 444 456.944444444444456.944\,444\,444\,444\,456.944 444 444 444 45% 66.964 285 714 285 7466.9642857142857466.964\,285\,714\,285\,7466.964 285 714 285 74% 22.222 222 222 222 2422.2222222222222422.222\,222\,$2 222 2$ 24% 52.605 459 057 071 9552.6054590570719552.605\,459\,057\,071\,9552.605 459 057 071 95% 23.333 333 333 333 3223.3333333333333223.333\,333\,$3 333 3$ 32% 46.571 428 571 428 55546.57142857142855546.571\,428\,571\,428\,55546.571 428 571 428 555% 58.297 872 340 425 5158.2978723404255158.297\,872\,340\,425\,5158.297 872 340 425 51% 37.953 091 684 434 96637.95309168443496637.953\,091\,684\,434\,96637.953 091 684 434 966% 51.094 890 510 948 951.094890510948951.094\,890\,510\,948\,951.094 890 510 948 9% 20.827 178 729 689 8120.8271787296898120.827\,178\,729\,689\,8120.827 178 729 689 81% 40.615 384 615 384 60640.61538461538460640.615\,384\,615\,384\,60640.615 384 615 384 606% 84.615 384 615 384 6184.6153846153846184.615\,384\,615\,384\,6184.615 384 615 384 61% 76.190 476 190 476 1876.1904761904761876.190\,476\,190\,476\,1876.190 476 190 476 18% 104.615 384 615 384 63104.61538461538463104.615\,384\,615\,384\,63104.615 384 615 384 63% 87.234 042 553 191 587.234042553191587.234\,042\,553\,191\,587.234 042 553 191 5% 92.325 581 395 348 8592.3255813953488592.325\,581\,395\,348\,8592.325 581 395 348 85% 50.446 428 571 428 58450.44642857142858450.446\,428\,571\,428\,58450.446 428 571 428 584%
Gemma BigRec 0.006 831 489 915 419 648 60.00683148991541964860.006\,831\,489\,915\,419\,648\,60.006 831 489 915 419 648 6 0.004 821 452 063 448 518 50.00482145206344851850.004\,821\,452\,063\,448\,518\,50.004 821 452 063 448 518 5 0.010 084 580 351 333 7670.0100845803513337670.010\,084\,580\,351\,333\,7670.010 084 580 351 333 767 0.005 833 862 406 063 7350.0058338624060637350.005\,833\,862\,406\,063\,7350.005 833 862 406 063 735 0.013 012 361 743 656 4730.0130123617436564730.013\,012\,361\,743\,656\,4730.013 012 361 743 656 473 0.006 600 741 721 637 5950.0066007417216375950.006\,600\,741\,721\,637\,5950.006 600 741 721 637 595 0.002 976 190 476 190 4760.0029761904761904760.002\,976\,190\,476\,190\,4760.002 976 190 476 190 476 0.002 976 190 476 190 476 50.00297619047619047650.002\,976\,190\,476\,190\,476\,50.002 976 190 476 190 476 5 0.005 208 333 333 333 3330.0052083333333333330.005\,208\,333\,333\,333\,3330.005 208 333 333 333 333 0.003 687 300 217 994 460 30.00368730021799446030.003\,687\,300\,217\,994\,460\,30.003 687 300 217 994 460 3 0.011 904 761 904 761 9040.0119047619047619040.011\,904\,761\,904\,761\,9040.011 904 761 904 761 904 0.005 312 154 716 450 0710.0053121547164500710.005\,312\,154\,716\,450\,0710.005 312 154 716 450 071 0.015 595 463 137 996 2190.0155954631379962190.015\,595\,463\,137\,996\,2190.015 595 463 137 996 219 0.010 489 160 983 015 6710.0104891609830156710.010\,489\,160\,983\,015\,6710.010 489 160 983 015 671 0.025 992 438 563 327 030.025992438563327030.025\,992\,438\,563\,327\,030.025 992 438 563 327 03 0.013 832 617 043 657 3280.0138326170436573280.013\,832\,617\,043\,657\,3280.013 832 617 043 657 328 0.043 005 671 077 504 730.043005671077504730.043\,005\,671\,077\,504\,730.043 005 671 077 504 73 0.018 162 307 633 210 0840.0181623076332100840.018\,162\,307\,633\,210\,0840.018 162 307 633 210 084
$D^{3}$ 0.007 161 458 333 333 3330.0071614583333333330.007\,161\,458\,333\,333\,3330.007 161 458 333 333 333 0.003 762 963 854 070 603 50.00376296385407060350.003\,762\,963\,854\,070\,603\,50.003 762 963 854 070 603 5 0.020 182 291 666 666 6680.0201822916666666680.020\,182\,291\,666\,666\,6680.020 182 291 666 666 668 0.007 995 190 196 701 8170.0079951901967018170.007\,995\,190\,196\,701\,8170.007 995 190 196 701 817 0.033 854 166 666 666 6640.0338541666666666640.033\,854\,166\,666\,666\,6640.033 854 166 666 666 664 0.011 421 699 038 261 1520.0114216990382611520.011\,421\,699\,038\,261\,1520.011 421 699 038 261 152 0.021 577 380 952 380 9520.0215773809523809520.021\,577\,380\,952\,380\,9520.021 577 380 952 380 952 0.012 923 434 799 794 6520.0129234347997946520.012\,923\,434\,799\,794\,6520.012 923 434 799 794 652 0.032 738 095 238 095 240.032738095238095240.032\,738\,095\,238\,095\,240.032 738 095 238 095 24 0.016 384 061 220 719 0350.0163840612207190350.016\,384\,061\,220\,719\,0350.016 384 061 220 719 035 0.044 642 857 142 857 1440.0446428571428571440.044\,642\,857\,142\,857\,1440.044 642 857 142 857 144 0.019 403 363 128 620 0520.0194033631286200520.019\,403\,363\,128\,620\,0520.019 403 363 128 620 052 0.011 657 214 870 825 4580.0116572148708254580.011\,657\,214\,870\,825\,4580.011 657 214 870 825 458 0.006 789 634 641 773 0710.0067896346417730710.006\,789\,634\,641\,773\,0710.006 789 634 641 773 071 0.020 951 480 781 348 4560.0209514807813484560.020\,951\,480\,781\,348\,4560.020 951 480 781 348 456 0.014 083 412 140 940 2040.0140834121409402040.014\,083\,412\,140\,940\,2040.014 083 412 140 940 204 0.037 807 183 364 839 320.037807183364839320.037\,807\,183\,364\,839\,320.037 807 183 364 839 32 0.022 433 535 336 669 4640.0224335353366694640.022\,433\,535\,336\,669\,4640.022 433 535 336 669 464
LangPTune 0.013 012 3620.0130123620.013\,012\,3620.013 012 362 0.007 850 0180.0078500180.007\,850\,0180.007 850 018 0.022 121 014 964 2160.0221210149642160.022\,121\,014\,964\,2160.022 121 014 964 216 0.010 718 4340.0107184340.010\,718\,4340.010 718 434 0.040 338 321 405 3350.0403383214053350.040\,338\,321\,405\,3350.040 338 321 405 335 0.015 231 8650.0152318650.015\,231\,8650.015 231 865 0.034 970 238 095 238 0960.0349702380952380960.034\,970\,238\,095\,238\,0960.034 970 238 095 238 096 0.023 467 457 188 027 2470.0234674571880272470.023\,467\,457\,188\,027\,2470.023 467 457 188 027 247 0.046 8750.0468750.046\,8750.046 875 0.027 365 138 310 761 6770.0273651383107616770.027\,365\,138\,310\,761\,6770.027 365 138 310 761 677 0.067 708 333 333 333 330.067708333333333330.067\,708\,333\,333\,333\,330.067 708 333 333 333 33 0.032 544 824 028 653 760.032544824028653760.032\,544\,824\,028\,653\,760.032 544 824 028 653 76 0.006 770 60.00677060.006\,770\,60.006 770 6 0.005 332 20.00533220.005\,332\,20.005 332 2 0.011 9670.0119670.011\,9670.011 967 0.005 852 70.00585270.005\,852\,70.005 852 7 0.019 5240.0195240.019\,5240.019 524 0.009 404 90.00940490.009\,404\,90.009 404 9
$\text{R}^{2}\text{ec}$ 0.026 350.026350.026\,350.026 35* 0.016 0810.0160810.016\,0810.016 081* 0.039 6880.0396880.039\,6880.039 688* 0.020 3050.0203050.020\,3050.020 305* 0.061 4830.0614830.061\,4830.061 483* 0.025 7280.0257280.025\,7280.025 728* 0.057 2920.0572920.057\,2920.057 292* 0.039 7510.0397510.039\,7510.039 751* 0.080 3570.0803570.080\,3570.080 357* 0.047 1540.0471540.047\,1540.047 154* 0.104 170.104170.104\,170.104 17* 0.052 6720.0526720.052\,6720.052 672* 0.032 593 292 394 898 440.032593292394898440.032\,593\,292\,394\,898\,440.032 593 292 394 898 44* 0.020 513 333 770 876 2880.0205133337708762880.020\,513\,333\,770\,876\,2880.020 513 333 770 876 288* 0.053 062 509 840 969 9260.0530625098409699260.053\,062\,509\,840\,969\,9260.053 062 509 840 969 926* 0.027 052 420 548 539 7840.0270524205485397840.027\,052\,420\,548\,539\,7840.027 052 420 548 539 784* 0.083 451 424 972 445 290.083451424972445290.083\,451\,424\,972\,445\,290.083 451 424 972 445 29* 0.034 727 909 699 741 6760.0347279096997416760.034\,727\,909\,699\,741\,6760.034 727 909 699 741 676*
% Improve. 50.857 142 8550.8571428550.857\,142\,8550.857 142 85% $11.8055$% 77.232 142 85777.23214285777.232\,142\,85777.232 142 857% 25.308 64225.30864225.308\,64225.308 642% $52.61$% 22.375 124 99422.37512499422.375\,124\,99422.375 124 994% 63.714 285 714 285 6963.7142857142856963.714\,285\,714\,285\,6963.714 285 714 285 69% 69.361 702 127 659 5869.3617021276595869.361\,702\,127\,659\,5869.361 702 127 659 58% 71.428 571 428 571 4471.4285714285714471.428\,571\,428\,571\,4471.428 571 428 571 44% 72.262 773 722 627 $\,722\,627\,73$ 722 627 73% 53.914 327 917 282 1453.9143279172821453.914\,327\,917\,282\,1453.914 327 917 282 14% 62.153 846 153 846 1462.1538461538461462.153\,846\,153\,846\,1462.153 846 153 846 14% 108.974 358 974 358 96108.97435897435896108.974\,358\,974\,358\,96108.974 358 974 358 96% 95.238 095 238 095 2395.2380952380952395.238\,095\,238\,095\,2395.238 095 238 095 23% 104.230 769 230 769 24104.23076923076924104.230\,769\,230\,769\,24104.230 769 230 769 24% 92.198 581 560 283 6992.1985815602836992.198\,581\,560\,283\,6992.198 581 560 283 69% 94.186 046 511 627 9494.1860465116279494.186\,046\,511\,627\,9494.186 046 511 627 94% 54.910 714 285 714 2954.9107142857142954.910\,714\,285\,714\,2954.910 714 285 714 29%
To validate the effectiveness of the proposed $\text{R}^{2}\text{ec}$, we showed the overall performance of baselines and our $\text{R}^{2}\text{ec}$ in Table 1. By analyzing the results, we gained the following findings. 1) Overall, $\text{R}^{2}\text{ec}$ consistently outperforms every competing baseline, underscoring the value of jointly optimizing reasoning and recommendation. 2) Traditional methods perform well on the Instruments dataset but struggle on CDs and Games, revealing their limited generality. LangPTune frequently ranks second, which validates the benefit of integrating explicit reasoning into the recommendation pipeline. Finally, large recommenders generally outperform traditional approaches, notably secure secondary position on the Games dataset. We attribute this advantage to their larger model scale and semantic understanding capabilities. 3) Comparing two backbones, Gemma consistently outperforms its larger counterpart, achieving up to 2× gains for $D^{3}$ – suggesting that Gemma may generally deliver stronger recommendation performance despite its smaller parameter count (2B vs. 3B).
5.3 Ablation Study
Table 2: Ablation study on key components of $\text{R}^{2}\text{ec}$.
Instruments CDs and Vinyl Video Games
Method H@5 N@5 H@10 N@10 H@20 N@20 H@5 N@5 H@10 N@10 H@20 N@20 H@5 N@5 H@10 N@10 H@20 N@20
w/o Reasoning 0.017 5670.0175670.017\,5670.017 567 0.012 0610.0120610.012\,0610.012 061 0.029 6030.0296030.029\,6030.029 603 0.015 321 020 0110.0153210200110.015\,321\,020\,0110.015 321 020 011 0.051 0740.0510740.051\,0740.051 074 0.020 0110.0200110.020\,0110.020 011 0.046 8750.0468750.046\,8750.046 875 0.032 141 739 768 641 3340.0321417397686413340.032\,141\,739\,768\,641\,3340.032 141 739 768 641 334 0.069 196 428 571 428 580.069196428571428580.069\,196\,428\,571\,428\,580.069 196 428 571 428 58 0.039 296 044 879 371 210.039296044879371210.039\,296\,044\,879\,371\,210.039 296 044 879 371 21 0.094 494 047 619 047 620.094494047619047620.094\,494\,047\,619\,047\,620.094 494 047 619 047 62 0.045 616 438 834 085 350.045616438834085350.045\,616\,438\,834\,085\,350.045 616 438 834 085 35 0.027 712 171 311 604 4720.0277121713116044720.027\,712\,171\,311\,604\,4720.027 712 171 311 604 472 0.017 357 424 748 561 4170.0173574247485614170.017\,357\,424\,748\,561\,4170.017 357 424 748 561 417 0.044 087 545 268 461 660.044087545268461660.044\,087\,545\,268\,461\,660.044 087 545 268 461 66 0.022 660 610 718 420 0330.0226606107184200330.022\,660\,610\,718\,420\,0330.022 660 610 718 420 033 0.074 791 371 437 568 90.07479137143756890.074\,791\,371\,437\,568\,90.074 791 371 437 568 9 0.030 337 079 249 673 180.030337079249673180.030\,337\,079\,249\,673\,180.030 337 079 249 673 18
w/o $R_{d}$ 0.019 8440.0198440.019\,8440.019 844 0.012 4110.0124110.012\,4110.012 411 0.033 8320.0338320.033\,8320.033 832 0.016 4140.0164140.016\,4140.016 414 0.055 9530.0559530.055\,9530.055 953 0.022 4250.0224250.022\,4250.022 425 0.052 0830.0520830.052\,0830.052 083 0.033 7820.0337820.033\,7820.033 782 0.076 6370.0766370.076\,6370.076 637 0.040 4270.0404270.040\,4270.040 427 0.097 4320.0974320.097\,4320.097 432 0.048 6430.0486430.048\,6430.048 643 0.030 2310.0302310.030\,2310.030 231 0.019 5720.0195720.019\,5720.019 572 0.048 6540.0486540.048\,6540.048 654 0.025 360.025360.025\,360.025 36 0.079 830.079830.079\,830.079 83 0.033 2230.0332230.033\,2230.033 223
w/o $R_{c}$ 0.024 3980.0243980.024\,3980.024 398 0.016 0360.0160360.016\,0360.016 036 0.039 3620.0393620.039\,3620.039 362 0.020 8190.0208190.020\,8190.020 819 0.060 5070.0605070.060\,5070.060 507 0.025 8470.0258470.025\,8470.025 847 0.054 3150.0543150.054\,3150.054 315 0.038 2270.0382270.038\,2270.038 227 0.077 3810.0773810.077\,3810.077 381 0.045 5870.0455870.045\,5870.045 587 0.101 190.101190.101\,190.101 19 0.051 4650.0514650.051\,4650.051 465 0.031 6490.0316490.031\,6490.031 649 0.020 2010.0202010.020\,2010.020 201 0.053 3770.0533770.053\,3770.053 377 0.026 3580.0263580.026\,3580.026 358 $0.0814$ 0.035 5220.0355220.035\,5220.035 522
$\text{R}^{2}\text{ec}$ 0.026 350.026350.026\,350.026 35 0.016 0810.0160810.016\,0810.016 081 0.039 6880.0396880.039\,6880.039 688 0.020 3050.0203050.020\,3050.020 305 0.061 4830.0614830.061\,4830.061 483 0.025 7280.0257280.025\,7280.025 728 0.058 780.058780.058\,780.058 78 0.038 8350.0388350.038\,8350.038 835 0.080 3570.0803570.080\,3570.080 357 0.045 6550.0456550.045\,6550.045 655 0.108 630.108630.108\,630.108 63 0.052 5010.0525010.052\,5010.052 501 0.032 593 292 394 898 440.032593292394898440.032\,593\,292\,394\,898\,440.032 593 292 394 898 44 0.020 513 333 770 876 2880.0205133337708762880.020\,513\,333\,770\,876\,2880.020 513 333 770 876 288 0.053 062 509 840 969 9260.0530625098409699260.053\,062\,509\,840\,969\,9260.053 062 509 840 969 926 0.027 110.027110.027\,110.027 11 0.085 3410.0853410.085\,3410.085 341 0.036 2930.0362930.036\,2930.036 293
We conducted ablation studies by evaluating the following variants: 1) “w/o Reasoning”. We eliminated reasoning tokens from prompts and trained exclusively with in-batch contrastive loss; 2) “w/o $R_{c}$”. We keep only the discrete ranking reward $R_{d}$ and removed the continuous similarity reward $R_{c}$ in Eqn.(5); and 3) “w/o $R_{d}$”. We removed the discrete ranking reward $R_{d}$ from Eqn.(5). The results are summarized in Table 2, and several observations stand out.
1) It is found that $\text{R}^{2}\text{ec}$ achieves an average improvement of roughly 15% across all metrics compared to w/o Reasoning. These substantial gains demonstrate that our designed optimization have enabled $\text{R}^{2}\text{ec}$ to better leverage test-time scaling to deliver significantly stronger recommendation performance. 2) As revealed that w/o $R_{c}$ (using only $R_{d}$) consistently outperforms w/o $R_{d}$ (using only $R_{c}$), this indicates that adopting reward signal that directly reflects evaluation result is crucial for training, while the continuous reward $R_{c}$, despite offering finer granularity, fails to provide meaningful distinctions and instead introduces noise that leads to suboptimal performance. 3) By fusing $R_{d}$ with a small weight on $R_{c}$, our approach preserves the task alignment of ranking rewards while benefiting from the supplementary signal of the continuous term. As a result, $\text{R}^{2}\text{ec}$ achieves optimal performance on nearly all metrics.
5.4 In-depth Analysis
(a) Train Reward (CDs)
(b) Val Reward (CDs)
(c) Val Length (CDs)
(d) Train Reward (Instruments)
(e) Val Reward (Instruments)
(f) Val Length (Instruments)
Figure 2: Analysis on advantage estimation methods, RLOO and GRPO, across two datasets. “Train Reward” and “Val Reward” indicate the variation in rewards on the training set and validation set, respectively. “Val Length” represents the variation in reasoning length on the validation set.
5.4.1 Analysis on Advantage Estimation
Accurate advantage estimation is crucial for reducing variance and improving sample efficiency in policy-gradient RL [26, 23, 34]. We therefore evaluated two estimators, GRPO [27] and RLOO [26], in our training pipeline. The variations of the training reward, validation reward, and reasoning length across the training steps are summarized in Figure 2, and we had the following observations.
1) First, as shown in Figure 2(a) and Figure 2(d), both RLOO and GRPO exhibit high-variance training curves. This is due to the inherent nature of recommendation environments, which produce highly varied reward magnitudes—some sessions result in high rankings, while others yield very low ones. 2) Second, as Figure 2(b) and Figure 2(e) illustrate, GRPO demonstrates faster learning in the initial training steps and consistently outperforms RLOO in terms of validation reward, whereas RLOO progresses more steadily. This divergence stems from the GRPO’s unit-variance normalization magnifies rewards into larger gradients that accelerate early learning. 3) Besides, as shown in Figure 2(c) and Figure 2(f), GRPO’s reasoning length gradually increases as training progresses, which is consistent with the phenomenon observed in LLM reasoning training [27], while RLOO appears to maintain a certain level of stability. This is because RLOO-driven optimization does not encourage improved reasoning with the low reward magnitudes in our task.
5.4.2 Analysis on Trajectory Sampling
For $\text{R}^{2}\text{ec}$’s training, we performed trajectory sampling via the temperature $\tau$ and top-$K$ sampling, which influence the stochasticity and diversity of the generated samples. To quantify their impact, we varied $\tau$ and $K$, and the results are presented in Figure 3. 1) It is observed that increasing the temperature produces longer resoning and boosts recommendation performance, i.e., , NDCG@5. A higher temperature introduces greater sampling entropy, allowing the model to explore a wider range of reasoning trajectories. 2) Conversely, it is found that increasing top-$K$ actually shortens the reasoning length and generally leads to a decline in recommendation performance. This is because a larger top-$K$ enlarges the candidate token set, which counterintuitively mitigates length hacking yet excessively large $K$ reintroduces noisy and low-quality samples.
Figure 3: Analysis on trajectory sampling and group size over the CDs dataset. (3(a)) and (3(b)) show the impact of temperature and top-$K$ sampling on performance and reasoning length, respectively. (3(c)) and (3(d)) present the effect of group size on NDCG@5 and Hit Rate@5, respectively.
5.4.3 Analysis on Group Size
During training, we sampled a group of trajectories to estimate their advantages. Therefore, we conducted experiments to analyze how varying group sizes impact performance. The results are outlined in Figure 3, and we gained several key findings as follows: 1) It is observed that performance improves for both backbones as group size increases, but the rate of improvement gradually slows down. While a larger group size generally leads to more explored paths, it also raises the training cost. These results suggest that selecting a group size of 6 or 8 is sufficient, and further increasing the group size is unnecessary. 2) Comparing the two backbones, we found a difference in sensitivity to group size. Qwen’s performance at a group size of 1 lags significantly behind Gemma’s, but it improves rapidly as the group size increases. Gemma performs well even with smaller groups, likely because its pretraining exposed it to a broader range of reasoning scenarios, thereby equipping it with stronger initial reasoning-for-recommendation capabilities.
5.4.4 More Analysis
Due to the scope limitation, we detail more analysis on the case study in Appendix H, and the analysis on the efficiency in Appendix G, respectively. We also discuss the limitations in Appendix F.
6 Conclusion and Future Work
[…]
Appendix A Dataset
Following the temporal-truncation protocol introduced by [3, 2], we construct three subsets—CDs and Vinyl (CDs), Video Games (Games), and Musical Instruments (Instruments) – from the lastest public Amazon review dataset [^1] https://amazon-reviews-2023.github.io/index.html. spanning from May 1996 to October 2023. For each domain, we begin with the most recent year of interactions (October 2022 – October 2023) and, if the resulting number of valid items is insufficient, iteratively roll the time window backward month-by-month until 10k items are obtained. We omit the 5-core filter so as to retain the nature behaviour characteristic of recommendation scenarios. Each user’s interaction history is then chronologically sorted and truncated to the latest 20 actions, yielding fixed-length sequences for all subsequent modelling and evaluation. Finally the dataset is split with 80% training 10% validation 10% test. The resulting statistics are listed in Table 3.
Appendix B Baselines
  • GRU4Rec [38] utilizes the GRU (Gated Recurrent Unit) architecture to model sequences.
  • Caser [39] employs both horizontal and vertical convolutional operations to enhance the capture of high-order interactions within item sequences, improving recommendation accuracy.
  • SASRec[40] incorporates a multi-head self-attention mechanism in its self-attentive sequential recommendation model, facilitating the modeling of intricate sequential data patterns.
  • TIGER [8] learns extra tokens from item features to present items, and then converts the user sequence into the sequence of the new item token for next-item generation.
  • BigRec [2] utilizes the item title to present the user sequence for recommendation generation.
  • $D^{3}$ [3] proposed to address amplification bias and homogeneity issues in LLM recommenders that generate item title generation.
  • LangPTune [15] utilized LLM as a user profile generator, and proposed an end-to-end training pipeline that optimizes LLM-generated user profiles for the recommendation objective.
Table 3: Dataset Statistics.
Dataset Users Items Density Interactions Train Val Test
Video Games 29,230 10,144 0.031% 63,502 50,801 6,350 6,351
CDs and Vinyl 7,701 12,024 0.023% 13,435 10,748 1,343 1,344
Musical Instruments 15,656 10,320 0.031% 34,373 27,498 3,437 3,438
Appendix C Implementation Details
  • Hardware: 4x NVIDIA 3090 (24GB) GPUs
  • Framework: PyTorch 2.5.0, Transformers 4.48.1, DeepSpeed 0.15.7
Traditional Recommenders.
We train all non-LLM baselines with the Cross-Entropy loss and the AdamW optimiser (learning-rate $1\times 10^{-4}$); the batch size is 256. TIGER [8] adopts T5 as its backbone without further architectural changes.
LLM-based Methods.
[…]
$\text{R}^{2}\text{ec}$ Optimization.
The learning rate is set as $1\text{e}{-5}$ with a batch size of 24, and apply linear warm-up for the first 32 steps. For trajectory generation, we set the group size $G$ as four, utilized vLLM [^2] https://docs.vllm.ai/en/latest/ (tensor-parallelism = 1, target GPU utilisation = 95 %) for efficient generation. vLLM reserves one GPU for inference and leaves three for training. Sampling uses temperature = 1.5 and top-$K$ = 200. Policy optimization follows the clipping ratio $\epsilon=0.2$ and omits KL regularization. Rewards are computed with $\text{NDCG}@1000$ and $\beta=0.05$. Unless stated otherwise, these settings are used throughout.
Appendix D Overall Pipeline
Training
To instantiate $\text{R}^{2}\text{ec}$, given the dataset $D$, the initial policy $\pi_{\theta}$, and item embedding cache $\mathbf{H}_{\mathcal{V}}$, we train the model with detailed training process illustrated in Algorithm 1.
Algorithm 1 Training Process
1:Dataset $\mathcal{D}$, initial policy $\pi_{\theta}$, embedding function $f_{\theta}$, item embedding table $\mathbf{H}_{\mathcal{V}}$
2:Optimized policy model $\pi_{\theta}$
3:for step $=1$ to $N$ do
4:     if step % $T_{\text{refresh}}==0$ then
5:         Refresh item embedding: $\mathbf{H}_{\mathcal{V}}[v]\leftarrow f_{\theta}({x}_{v}),\quad\forall v\in\mathcal{V}$
6:     end if
7:     Sample a training batch =$\mathcal{B}=\{(u,v^{+})\}\sim\mathcal{D}$
8:     Encode target item prompts and update embedding table: $\mathbf{H}_{\mathcal{V}}[v^{+}]\leftarrow f_{\theta}({x}_{v^{+}})\quad\forall(u,v^{+})\in\mathcal{B}$
9:     for all $(u,v^{+})$ in $\mathcal{B}$ do
10:         Generate $G$ trajectory: $\{[{o}_{1},v^{+}]...,[{o}_{G},v^{+}]\}\sim\pi_{\theta_{\text{old}}}(\cdot|{x}_{u})$
11:         Compute reward for each trajectory using Eq. (5)
12:         Compute advantage for each trajectory using Eq. (2)
13:     end for
14:     Update policy parameters $\theta$ via loss in Eq. (8)
15:     Update old policy: $\theta_{\text{old}}\leftarrow\theta$
16:end for
Inference
For inference, we first pass every item prompt $x_{v}$ through the trained model once to obtain all item embeddings $\mathbf{H}_{\mathcal{V}}$. At inference time, the model greedily generates a deterministic reasoning $o$ for a user prompt $x_{u}$, the last hidden state $\bm{h}_{T}$ goes through the embedding table, scoring each candidate via the inner product $s(v)=\bm{h}_{T}^{\top}\mathbf{H}_{\mathcal{V}}[v]$. The top-$K$ items can then be recommended by sorting these scores.
Appendix E Prompt
User Prompt Analyze in depth and finally recommend next {category} I might purchase inside <answer> and </answer>  . For example, <answer> a product </answer>  . Below is my historical {category} purchases and ratings (out of 5): {for hist in purchase_histories %} {{hist.time_delta} ago: [{hist.item_title}] ({hist.rating}) %}
Item Prompt Summarize key attributes of the following {category} inside <answer> and </answer>  : {for key, attr in item.meta %} {{key}: {attr} %}
Figure 4: Prompt templates for user interaction history and item metadata. The User Prompt encodes a user’s past purchases as a sequence of item titles, relative timestamps (e.g., “2hrs”, “4d”), and explicit ratings (in $[1,5]$), followed by an instruction to analyze and recommend the next item within the span of </answer>  and </answer>  . The Item Prompt summarizes structured item attributes (e.g., brand, type, features) with the same format requirement.
As illustrated in Figure 4, user interaction histories and item metadata are serialized into token sequences via the given prompt templates. To signal the boundary between reasoning and recommendation, we introduce control symbol <answer> , which triggers a shift from language modeling head to recommendation head.
Appendix F Limitations
Our work acknowledges two primary limitations. First, introducing explicit reasoning generation inevitably increases inference latency and reduces efficiency due to additional autoregressive decoding steps; nevertheless, exploring the potential of reasoning within recommendation is valuable, and our experiments have empirically confirmed its effectiveness in improving recommendation accuracy. Second, constrained by computational resources, we employed parameter-efficient tuning (LoRA) rather than full-parameter fine-tuning, thus not fully demonstrating the potentially superior performance achievable through comprehensive optimization.
Appendix G Analysis on Inference Efficiency
Table 4: Average per-request inference latency (mean) of our unified model $\text{R}^{2}\text{ec}$ versus the two-stage LangPtune pipeline, measured across varying batch sizes.
batch size $\text{R}^{2}\text{ec}$ LangPTune
1 16725.54 $\pm$ 2636.84 19030.95 $\pm$ 2753.81
4 3980.77 $\pm$ 77.13 4625.045 $\pm$ 67.00
8 2220.88 $\pm$ 32.00 2375.64 $\pm$ 69.476723
To quantify the efficiency gains of our unified model $\text{R}^{2}\text{ec}$ over the two-stage conventional pipeline, we benchmark its inference latency against LangPtune [15]. Inferences were performed on one NVIDIA RTX 3090 GPU using identical input prompts, maximum token lengths, and candidate sets. We executed 100 queries at batch sizes of 1, 4, and 8. For each batch, per-request latency was calculated by dividing the total duration by the batch size. The mean $\pm$ standard deviation for each configuration is reported in Table 4. $\text{R}^{2}\text{ec}$ achieves approximately 10% reduction in average per-request inference time, demonstrating the superior efficiency of our model architecture compared to the conventional two-stage approach. Notably, the efficiency gap narrows as batch size increases, since parallelization spreads computation costs over more requests. In addition, the low standard deviations demonstrate that $\text{R}^{2}\text{ec}$’s inference speed is reliably stable.
Appendix H Analysis on Case Study
To illustrate domain-specific and backbone-specific reasoning behaviors, we analyze representative cases from the CDs and Video Games datasets (Figures 5,6,7,8) using Gemma and Qwen as backbones. 1) On the CDs dataset, both models tend to reason by discussing several potentially interesting items and explicitly enumerating their features and genres, before providing a final recommendation. This behavior highlights the models’ ability to leverage domain knowledge acquired during pretraining and reinforced through RecPO, for better recommendation. 2) On the Video Games dataset, the models shift towards inferring user intent from the purchase sequence, speculating on the likely category or intended use of the next purchase. This demonstrates a shift from feature-based enumeration to intent-driven reasoning as the domain changes. 3) Notably, Gemma generally produces shorter and more concise reasoning sequences compared to Qwen, which often generates more elaborate justifications. These patterns indicate that, while both models benefit from explicit reasoning, their generation style and reliance on pretrained knowledge can manifest differently across domains and backbone choices.
Case Study 1 (CD and Vinyl) User Analyze in depth and finally recommend next CD or vinyl I might purchase inside <answer>and </answer>. For example, <answer>a product</answer>. Below is my historical CDs and vinyl purchases and ratings (out of 5): 255d 13.2h ago: [Pure 60’s: The #1 Hits] (5.0) 255d 12.6h ago: [Seven] (4.0) 26.5min ago: [Comeblack] (4.0) 18.9min ago: [Hot Rocks 1964-1971] (5.0) 16.3min ago: [Return to Forever] (4.0) 9.4min ago: [Rock Believer] (5.0) 6.3min ago: [From Small Things: The Best of Dave Edmunds] (3.0) $\text{R}^{2}\text{ec}$ (Gemma2-2B) Based on your past purchases, it seems you enjoy a mix of classic rock, soul/funk, and a touch of progressive rock. You appreciate well-crafted albums with strong vocals and a focus on musicality. Here are some recommendations, keeping in mind your taste: CD/Vinyl: The Very Best of The Rolling Stones (CD/Vinyl): You’ve enjoyed classic rock, and this compilation offers a great introduction to the Stones’ iconic sound. It features their biggest hits and showcases their bluesy roots and rock ’n’ roll energy. The Dark Side of the Moon (Vinyl): This Pink Floyd masterpiece is a classic for a reason. It’s a complex and innovative album with a timeless sound. The vinyl format adds a special touch to the listening experience. Kind of Blue (Vinyl): Miles Davis’s iconic jazz album is a must-have for any music lover. It’s a beautiful and introspective album that showcases Davis’s genius as a musician. The vinyl format brings out the nuances of the music. The Miseducation of Lauryn Hill (CD/Vinyl): This album is a powerful blend of hip-hop, soul, and R&B. Lauryn Hill’s incredible vocals and the album’s socially conscious lyrics make it a truly special listening experience. Why these recommendations: Variety: These albums offer a mix of genres and styles, reflecting your diverse taste. Quality: They are all critically acclaimed and considered essential listening. Format: The vinyl format adds a unique and immersive experience, while the CD format offers convenience and portability. Ultimately, the best way to find your next favorite album is to explore and discover new music. I recommend<answer>
Figure 5: Case Study on CDs with Gemma as backbone.
Case Study 2 (CD and Vinyl) User Analyze in depth and finally recommend next CD or vinyl I might purchase inside <answer>and </answer>. For example, <answer>a product</answer>. Below is my historical CDs and vinyl purchases and ratings (out of 5): 97d 7.7h ago: [Live At The El Mocambo[2 CD]] (5.0) 97d 7.5h ago: [Radio Silence] (5.0) 2d 16.5h ago: [Little Box ’O’ Snakes The Sunburst Years 1978-1982] (5.0) 2d 16.0h ago: [Goats Head Soup Super Deluxe] (5.0) $\text{R}^{2}\text{ec}$ (Qwen2.5-3B) Given your history of purchasing and rating albums, it seems you have a strong preference for live recordings, classic rock, and deluxe editions. Based on this, here are some recommendations for CDs or vinyls that align with your tastes: 1. “The Wall” (Deluxe Edition)
Artist:
Pink Floyd
Format: CD or Vinyl
Description: This is a legendary album that has been reissued multiple times, including a deluxe edition that includes rare tracks, documentaries, and bonus material. It’s a must-have for any Pink Floyd fan.
Rating: Given your appreciation for deluxe editions and classic rock, this would likely receive a high rating.
2. “The Dark Side of the Moon” (Super Deluxe Edition)
Artist:
Pink Floyd
Format: CD or Vinyl
Description: Another iconic album that has been extensively reissued. The super deluxe edition includes a wealth of bonus content, making it a comprehensive collection for fans.
Rating: Similar to “Goats Head Soup,” this would also likely receive a high rating due to its extensive content and your appreciation for deluxe editions.
3. “Live at the Royal Albert Hall” by David Bowie
Artist:
David Bowie
Format: CD or Vinyl
Description: This live album captures one of David Bowie’s most famous performances, showcasing his versatility and the energy of his live shows. It’s a great addition to your collection if you enjoy live recordings.
Rating: Given your enjoyment of live recordings, this would likely be highly rated.
4. “The Complete Studio Albums” by Fleetwood Mac
Artist:
Fleetwood Mac
Format: CD or Vinyl
Description: This box set includes all of Fleetwood Mac’s studio albums, providing a comprehensive overview of their career. It’s a great way to explore their evolution as a band.
Rating: If you’re a fan of Fleetwood Mac, this would likely be highly rated due to the quality of their music and the comprehensive nature of the collection.
5. “The Best of the Beatles” (Deluxe Edition)
Artist:
The Beatles
Format: CD or Vinyl
Description: A compilation that includes some of the Beatles’ most popular songs, arranged in chronological order. The deluxe edition often includes additional content like rare recordings and documentaries.
Rating: Given your appreciation for deluxe editions and classic rock, this would likely receive a high rating.
Conclusion:
Based on your past purchases and ratings, you might want to consider “The Wall” (Deluxe Edition), “The Dark Side of the Moon” (Super Deluxe Edition), “Live at the Royal Albert Hall” by David Bowie, “The Complete Studio Albums” by Fleetwood Mac, or “The Best of the Beatles” (Deluxe Edition). Each of these options offers a unique and comprehensive look into the artists’ work, which aligns with your tastes.
Figure 6: Case Study on CDs with Qwen as backbone.
Case Study 3 (Video Games) User Analyze in depth and finally recommend next video game I might purchase inside <answer>and </answer>. For example, <answer>a product</answer>. Below is my historical video games purchases and ratings (out of 5): 80d 1.6h ago: [For Apple Watch Fast Charger USB C, 2 in 1 Apple MFi Certified 3FT iWatch Charger Magnetic Charging Cable &PD USBC Fast Wall Charger Power Adapter for Apple Watch Series 6/SE/5/4/3/2/1&iPhone 14/13/12] (5.0) 76d 7.5h ago: [Kickstand Back Shell Holder Back Bracket for Nintendo Switch Console, Switch Replacement Kickstand Repair Kit with Screwdriver] (4.0) 45d 12.5h ago: [VOYEE Switch Controllers Wireless, Compatible with Switch Pro Controllers for Switch/Switch Lite/OLED/PC, One-Key Pairing Gamepad with 9-Color RGB LEDs 800mAh Battery 6-Axis Motion Wake Up Function] (5.0) 40d 23.5h ago: [Gvyugke 2.4GHz Wireless Gaming Headset for PS4, PS5, PC, Nintendo Switch, Bluetooth 5.2 Gaming Headphones with Mic for Mobile Device, Noise Canceling, Bass Surround, 50mm Driver, 40H Battery] (5.0) $\text{R}^{2}\text{ec}$ (Gemma2-2B) It seems you enjoy tech accessories and gaming peripherals! Based on your recent purchases, you likely appreciate: Functionality: You’re drawn to products that enhance your gaming experience, like the charging cable and the wireless gaming headset. Convenience: You value products that simplify your life, like the kickstand for your Switch. Quality:You prioritize products that are well-made and reliable, evidenced by your high ratings for the charging cable and the gaming headset. Considering these preferences, here are some recommendations for your next video game purchase: <answer>
Figure 7: Case Study on Games with Gemma as backbone.
Case Study 4 (Video Games) User Analyze in depth and finally recommend next video game I might purchase inside <answer>and </answer>. For example, <answer>a product</answer>. Below is my historical video games purchases and ratings (out of 5): 79d 15.5h ago: [BUSQUEDA Pro Head Strap with Battery for Quest 2, Magnetic Connection 6000mAh Replaceable VR Power, Lightweight & Balance Elite Strap Meta Accessories for Enhanced Support and Comfort in VR] (5.0) 69d 14.2h ago: [iTwano Stainless VR Stand for Quest 2 Pro, Desk Mount VR Holder and Headset Stand Compatible with Quest Pro 2 Rift S] (5.0) 59d 18.9h ago: [ROCCAT Vulcan II Max Optical-Mechanical PC Gaming Keyboard with Customizable RGB Illuminated Keys and Palm Rest, Titan II Smooth Linear Switches, Aluminum Plate, 100M Keystroke Durability Black] (5.0) 57d 21.9h ago: [ZIYOU LANG K3 Mechanical Gaming Keyboard 98 Keys with Number pad Chroma RGB Illuminated Water-Resistant Anti-Ghosting NKRO Retro Metal Panel Type-C for PC/Mac Gamer Typist Optical Clicky Blue Switch] (5.0) 57d 21.8h ago: [Corsair K60 PRO TKL Wired Optical-Mechanical OPX Linear Switch Gaming Keyboard with 8000Hz Polling Rate - Black] (4.0) 52d 16.8h ago: [IROK FE87/104 RGB Mechanical Keyboard, Hot Swappable Gaming Keyboard, Customizable Backlit, Magnet Upper Cover Type-C Wired Keyboard for Mac Windows-White/Brown Switch] (5.0) $\text{R}^{2}\text{ec}$ (Qwen2.5-3B) Given your purchase history, it is clear that your interests focus primarily on VR setup accessories, gaming peripherals (particularly high-quality keyboards), and devices that enhance comfort and customizability during play, rather than on specific video games themselves. Based on the trends observed in your history: VR Content: Multiple purchases of Quest 2/Pro accessories indicate a strong inclination towards immersive VR experiences. Comfort & Ergonomics: Investment in premium straps and ergonomic keyboards suggests that gaming comfort is a high priority. Customizability and RGB Lighting: The frequent selection of customizable RGB peripherals demonstrates your preference for visually dynamic and personalized setups. With these points from your purchases, let’s recommend one of the following types of next-gen games featuring improved immersive experiences for VR: <answer>
Figure 8: Case Study on Games with Qwen as backbone.